demolisher1224:

evaluate the function for x=2 f(x)=7x-5 no answers so u have to give me the solution

demolisher1224:

@563blackghost @theDeviliscoming

563blackghost:

Plug in x. We know that x=2, so plug in 2. \(\large\bf{f(2)=7(2)-5}\)

demolisher1224:

ok thx ghost

563blackghost:

np

demolisher1224:

ok hod on one minute the answers are right here A.9 B.-9 C.19 D.-19

563blackghost:

Did you do the equation?

demolisher1224:

uh no

demolisher1224:

ok its -9

563blackghost:

not quite... \(\large\bf{7(2) = 14}\) \(\large\bf{14-5}\)

demolisher1224:

9

563blackghost:

that's your answer..

ronndjaia:

replace x = 2 into f(x)=7x-5 f(2)=7(2)-5 = 14 -5 = 9 answer is 9